OpenStudy (anonymous):

The Mills Lunch Shop prepares fresh take-out entrees each day. On Tuesday, 40 baked chicken dinners were prepared at a cost of $3.20 each. A 20% spoilage rate is anticipated. Mills Lunch Shop sells the dinners at an 80% markup based on cost. They decided to offer a $1.00-off coupon in a newspaper advertisement. What markdown percentage does this coupon represent

OpenStudy (asevilla5):

Cost Price of 40 chickens = 40 * 3.2 = 128 Effective Cost Price = spoilage + 128 = 20% of 128 + 128 = 1.20*128 = 153.6 Profit = 80% Profit of 40 chickens = 153.6* 80/100 = 122.8 S.P of 40 chickens= 122.8 + 153.6 = 276.48 S.P of 1 chicken = 276.48/40 = 6.912 markdown amount = $1 markdown percentage = (markdown amount/SP)*100 markdown percentage = (1/6.912)*100 markdown percentage = 14.5%